# Multiscale Quantile Regression with Local Error Control

### v2

@@ 2.2 The proposed MUSCLE
A reasonable candidate $g$ for the recovery of the quantile function $f$ in Model 1 should pass the multiscale tests on individual segments. Thus, we restrict ourselves to functions lying in the multiscale side-constraint [^1] In multiscale segmentation methods (e.g. Frick et al., , 2014; Li et al., , 2016; Pein et al., , 2017; Li et al., , 2019; Vanegas et al., , 2022), the side constraint takes the form of $T_{I}(Z,\theta)-q_{\alpha}(\lvert I\rvert)\leq 0$, whereas in (5) we use $\accentset{\circ}{I}$ instead of $I$. This subtle adjustment ensures the independence of test statistics on adjacent segments and allows improved bounds on the distributional behaviours. For instance, in later Theorem 2, with such an adjustment, we can improve the upper bound of probability $\mathbbm{P}(\widehat{K}\geq K+k)$ from $\alpha^{\lfloor{k/2}\rfloor}$ to $\alpha^{k}$ for all $k\in\mathbbm{N}$.
$\mathcal{C}_{k}\coloneqq\biggl{\{}g=\sum_{j=0}^{k}\theta_{j}\mathbbm{1}_{{I}_{j}}\;:\;\bigsqcup_{j=0}^{k}I_{j}=[0,1)\text{ and }T_{\accentset{\circ}{I}_{j}}(Z,\theta_{j})-q_{\alpha}(\lvert\accentset{\circ}{I}_{j}\rvert)\leq 0,\,j=0,\dots,k\biggr{\}},$ (5)
where $\accentset{\circ}{I}$ denotes the interior of interval $I$. Note that in this particular form of $\mathcal{C}_{k}$ we control the type I error locally on every segment $I_{i}$. Further, to avoid overfitting to the data, we minimise the complexity of candidates, which is measured by the number of change points, and define
@@
Note that $\tilde{f}=\sum_{i=1}^{n}Z_{i}\mathbbm{1}_{[x_{i},x_{i+1})}$ clearly lies in $\mathcal{C}_{n-1}$, so $\widehat{K}$ is finite and takes values in $\{0,1,\ldots,n-1\}$. Thus $\mathcal{C}_{\widehat{K}}\neq\emptyset$ and consequently a solution $\widehat{f}$ of (7) always exists. Owing to the specific structure of the optimisation problem, $\widehat{f}$ can be efficiently computed in the dynamic programming framework (see Section 4).
@@ 2.3 Comparison with global error control
As a simulation study, we consider the recovery of the teeth function that contains $K=80$ change points under the additive noise model (1), where the noises are scaled $t_{3}$-distributed with standard deviation equal to one. The simulation results are given in Figure 2 and Table 1. Clearly, MUSCLE ($\alpha=0.3$) recovers nicely the number and the locations of change points, while MQS ($\tilde{\alpha}=0.3$) severely underestimates the number of change points. To enhance the detection power of MQS, we adjust $\tilde{\alpha}=1-(1-0.3)^{81}\approx 0.999$. This adjustment significantly improves MQS in the estimation of $K$, bringing its performance closer to MUSCLE, but at the expense of nearly no statistical guarantee. Still, MQS with such a large $\tilde{\alpha}$ underperforms MUSCLE in estimating the change point locations as well as the whole signal. See Section 5.1 (and also Section 1.2) for further comparisons.
@@
Table 1: Comparison of MQS and MUSCLE in the setup of Figure 2. For each criterion, the average value and the standard deviation (in brackets), over 200 repetitions, are presented. Here $d_{H}$ is the Hausdorff distance between estimated and true change points, MISE the mean integrated squared error of $\widehat{f}$ and MIAE the mean integrated absolute error. Their formal definitions are given in (12), and in (24) and (25) in Appendix B, respectively.
@@
$\widehat{K}$ $80\;(0.4)$ $40\;(0.5)$ $77\;(1.6)$
$d_{H}$ $0.001\;(0.0030)$ $0.015\;(0.0037)$ $0.008\;(0.0006)$
MISE $0.1\;(0.041)$ $2.7\;(0.065)$ $1.8\;(0.108)$
MIAE $0.2\;(0.020)$ $1.4\;(0.008)$ $0.9\;(0.057)$
@@ 3 Statistical Properties
3.1 Consistency and localisation rates
@@
$F_{\min,k}=\underset{i:x_{i}\in I_{k}}{\min}F_{i},\quad\text{and}\quad F_{\max,k}=\underset{i:x_{i}\in I_{k}}{\max}F_{i}.$
For $k=1,\ldots,K$, let $\delta_{k}=\theta_{k}-\theta_{k-1}$ and refer to its absolute value as jump size. We use $\xi_{k}=\min\left\{\xi_{F_{\min,k},\beta}(\delta_{k}/2),\xi_{F_{\max,k},\beta}(-\delta_{k}/2)\right\}$ to characterise the quantile jump at $\tau_{k}$. In case of $F_{i}$ being continuously differentiable, $\xi_{k}$ is equivalent to $\lvert\delta_{k}\rvert$ up to a constant if $\lvert\delta_{k}\rvert$ is small, see Example 1. Further, we define
$\Lambda\coloneqq\underset{k=1,\dots,K}{\min}\lvert\tau_{k}-\tau_{k-1}\rvert,\quad\Theta\coloneqq\underset{k=1,\dots,K}{\min}\xi_{k},\quad\text{and}\quad\Omega\coloneqq\underset{k=1,\dots,K}{\min}\sqrt{\lambda_{k}}\xi_{k}.$ (8)
@@
    lim $\liminf\Lambda_{n}>0$ and $\sqrt{n}\Omega_{n}-\max_{m\leq n}q_{\alpha_{n}}(m)/\sqrt{2}\to\infty$.
@@
    lim $\liminf\Lambda_{n}=0$ and for some constant $c>2$,
    $\sqrt{n}\Omega_{n}\geq c\sqrt{-\log\Lambda_{n}},\text{ and }(c-2)\sqrt{-2\log\Lambda_{n}}-{\max_{m\leq n}q_{\alpha_{n}}(m)}\to\infty.$
@@
By Lemma 10 in Appendix A we have $\max_{m\leq n}q_{\alpha_{n}}(m)\lesssim\sqrt{-\log\alpha_{n}}.$ This allows us to select $\alpha_{n}$ approaching zero at certain explicit rates to satisfy the above conditions on $q_{\alpha_{n}}(m)$. We discuss the implication of Theorem 3 under conditions (i) and (ii), separately.
@@
    Note that $\liminf_{n}\Lambda_{n}>0$ implies that $\sup K_{n}<\infty$, that is, the number of change points stays bounded. In particular, if the quantile function $f=f_{n}$ in Model 1 stays the same for different $n$, MUSCLE with any zero sequence $\alpha_{n}\searrow 0$ estimates consistently the number of change points.
@@
    Consider the example of $f_{n}=\Delta_{n}\mathbbm{1}_{[0.5,0.5+\Lambda_{n})}$ with additive standard Gaussian noises and jump size $\Delta_{n}>0$, which corresponds to Model 1 with $\beta=0.5$. By simple computation we have
@@
Let $T$ and $\widehat{T}$ denote the sets of true change points and estimated change points, respectively. The localisation error between $T$ and $\widehat{T}$ is defined by
@@
[…]
Theorem 4 (Localisation rate in Hausdorff distance).
@@
$\mathbb{P}\left(d_{H}(T_{n},\widehat{T}_{n})\geq\varepsilon_{n}\right)\leq\alpha_{n}+4K_{n}\exp\left(-n\varepsilon_{n}\Theta_{n}^{2}\right)+\\ 4K_{n}\exp\left[-\Bigl{(}\sqrt{n\varepsilon_{n}}\Theta_{n}-\tfrac{\underset{m\leq n}{\max}q_{\alpha_{n}}(m)}{\sqrt{2}}-\sqrt{\log\tfrac{2e}{\varepsilon_{n}}}\Bigr{)}_{+}^{2}\right]$ (13)
@@
$\varepsilon_{n}=\frac{C\log n}{n\Theta_{n}^{2}}\quad\text{for some sufficiently large constant C.}$.
@@ 3.2 Control of false positives
[…]
@@ 4 Algorithm and Implementation
$\widehat{K}[j]\coloneqq\min\Bigg{\{}k\,:\underset{0\leq i\leq k}{\max}T_{\accentset{\circ}{I}_{i}}(Z,\theta_{i})-q_{\alpha}(|\accentset{\circ}{I}_{i}|)\leq 0,\text{ for some }f=\sum_{i=0}^{k}\theta_{i}\mathbbm{1}_{I_{i}}\text{ with }\bigsqcup_{i=0}^{k}I_{i}=\bigl{[}0,\tfrac{j}{n}\bigr{)}\Bigg{\}}.$
@@
$\mathcal{I}\left(I\right)=\begin{cases}1&\text{if there is a $\theta$ such that }T_{\accentset{\circ}{I}}(Z,\theta)\leq q_{\alpha}(\lvert\accentset{\circ}{I}\rvert),\\ 0&\text{otherwise.}\end{cases}$
@@
Finally, combining the wavelet tree and the dynamic programming with pruning techniques, we obtain Algorithm 1 (in Appendix C) for computing the global optimal solution to the non-convex optimisation problem defining MUSCLE.
@@
MUSCLE, defined as the global optimal solution to the optimisation problem in (5)–(7), can be computed by a pruned dynamic program with the wavelet tree data structure, which has a space complexity (i.e., amount of memory) of $O(n\log n)$ and a worst-case computational complexity (i.e., runtime) of $O(n^{4}\log n)$. Further, if we use the intervals of dyadic lengths in MUSCLE, the worst-case computational complexity is $O(n^{3}\log^{2}n)$.
@@ 5.1 Simulation studies
We investigate the empirical performance of MUSCLE and its splitting-merging variant, denoted by MUSCLE-S (see Remark 4), under various situations. To benchmark the performance, we include PELT (Killick et al., , 2012), WBS (Fryzlewicz, , 2014), SMUCE (Frick et al., , 2014), FDRSeg (Li et al., , 2016) and MQS (Vanegas et al., , 2022) in Gaussian models, and R-FPOP (Fearnhead and Rigaill, , 2019), KSD (Madrid Padilla et al., , 2021), MQS (Vanegas et al., , 2022), NOT-HT (Baranowski et al., , 2019, the heavy tailed version), ED-PELT (Haynes et al., , 2017), RNSP (Fryzlewicz, , 2024, the CUSUM version) in non-Gaussian models, as competitors. For implementation, we compute WBS from R package wbs, PELT from changepoints, SMUCE and FDRSeg from FDRSeg, MQS from mqs, NOT-HT from not, ED-PELT from changepoint.np, all available from CRAN. The implementations of R-FPOP, KSD and RNSP are available on Github (https://github.com/guillemr/robust-fpop, https://github.com/hernanmp/NWBS, https://github.com/pfryz/nsp). Default tuning parameters are used for PELT, R-FPOP, WBS, KSD and NOT-HT. For SMUCE, MQS, FDRSeg MUSCLE, MUSCLE-S and RNSP, we select $\alpha=0.3$. The length of pieces in MUSCLE-S is set to $300$. We use the max_sign_cusums criterion (see Sen and Srivastava, , 1975 and Fryzlewicz, , 2024) for estimating change points in RNSP.
As quantitative evaluation criteria of change points, we consider the estimated number $\widehat{K}$ of change points, the localisation error $d(\cdot;\cdot)$ in (11), the Hausdorff distance $d_{H}(\cdot,\cdot)$ in (12), and the FDR in (14). We also include the standard criteria of function estimation, consisting of the mean integrated square error (MISE) and the mean integrated absolute error (MIAE), formally defined in (24) and (25) in Appendix B. Additionally, as change point problems can be viewed as clustering problems, we incorporate the V-measure (Rosenberg and Hirschberg, , 2007) which ranges from 00 to $1$ with larger values indicating higher accuracy.
@@ 5.1.1 Changes in median
We consider the additive noise model in (1) with $\beta=0.5$, i.e., noises $\varepsilon_{i}$ are independent random variables with median zero. The test signal in the Gaussian scenario (E1) has $K=2$ change points at $986$ and $1016$ with sample size $n=2000$. The values between change points are $-4$, 00 and $4$. In other scenarios (E2)(E5) we use the blocks signal of length $n=2048$ (Donoho and Johnstone, , 1994). Its change points are 205, 267, 308, 472, 512, 820, 902, 1332, 1557, 1598, 1659 and the values between change points are 00, $14.64$, $-3.66$, $7.32$, $-7.32$, $10.98$, $-4.39$, $3.29$, $19.03$, $7.68$, $15.37$, 00. The noise distributions are specified as follows.
@@
    Normal distribution: $\varepsilon_{i}\sim\mathcal{N}(0,0.9)$.
@@
    Scaled $t$-distribution: $\varepsilon_{i}\sim 2^{-0.5}\cdot\bigl{(}8\cdot\mathbbm{1}_{\{1\leq i\leq 389\}}+0.5\cdot\mathbbm{1}_{\{390\leq i\leq 666\}}+4\cdot\mathbbm{1}_{\{667\leq i\leq 1445\}}+\mathbbm{1}_{\{1446\leq i\leq 2048\}}\bigr{)}\cdot t_{3}$
@@
    Scaled Cauchy distribution: $\varepsilon_{i}\sim\bigl{(}0.6\cdot\mathbbm{1}_{\{1\leq i\leq 389\}}+0.05\cdot\mathbbm{1}_{\{390\leq i\leq 666\}}+0.6\cdot\mathbbm{1}_{\{667\leq i\leq 1445\}}+0.2\cdot\mathbbm{1}_{\{1446\leq i\leq 2048\}}\bigr{)}\cdot{\rm Cauchy(0,1)}.$
@@
    Scaled centred $\chi^{2}$-distribution: $\varepsilon_{i}\sim 6^{-0.5}\cdot\bigl{(}6\cdot\mathbbm{1}_{\{1\leq i\leq 389\}}+0.5\cdot\mathbbm{1}_{\{390\leq i\leq 666\}}+6\cdot\mathbbm{1}_{\{667\leq i\leq 1445\}}+2\cdot\mathbbm{1}_{\{1446\leq i\leq 2048\}}\bigr{)}\cdot\bigl{(}\chi_{3}^{2}-\text{median}(\chi_{3}^{2})\bigr{)}.$
@@
[…]
In all scenarios, the simulations are repeated 200 times. We display simulation results in scenarios (E1)(E2) and (E5) in Figures 3, 4 and 5, respectively. Further detailed results as well as scenarios (E3) and (E4) are given in Appendix B (Figures 9, 10 and 2). In the Gaussian scenario (E1), FDRSeg and WBS overestimate the number of change points, leading to a relatively large FDR. In contrast, PELT, SMUCE and MUSCLE provide more accurate estimation of $K$. Despite MUSCLE having a slightly larger MISE compared to SMUCE and FDRSeg, it outperforms PELT and WBS. Overall, the performance of MUSCLE is comparable to the procedures tailored for the Gaussian model. In non-Gaussian scenarios (E2)(E5), there are 11 change points in median, and 14 change points in distribution. We compute localisation error, Hausdorff distance, FDR and V-measure with respect to the change points in median, and MISE and MIAE with respect to the median function. We discuss the performance of each method in detail as follows.
@@
  • KSD aims at the detection of distributional changes, but often fails to detect all of $14$ change points. The detected change points contain only a subset of change points in median, as revealed by Hausdorff distance and MISE (as well as localisation error, MIAE and V-measure in Table 2 in Appendix B).

### v3

@@ 2.2 The proposed MUSCLE
A reasonable candidate $g$ for the recovery of the quantile function $f$ in Model 1 should pass the multiscale tests on individual segments. Thus, we restrict ourselves to functions lying in the multiscale side-constraint [^1] In multiscale segmentation methods (e.g. Frick et al., , 2014; Li et al., , 2016; Pein et al., , 2017; Li et al., , 2019; Vanegas et al., , 2022), the side constraint takes the form of $T_{I}(Z,\theta)-q_{\alpha}(\lvert I\rvert)\leq 0$, whereas in (5) we use $\mathring{I}$ instead of $I$. This subtle adjustment ensures the independence of test statistics on adjacent segments and allows improved bounds on the distributional behaviors. For instance, in later Theorem 2, with such an adjustment, we can improve the upper bound of probability $\mathbbm{P}(\widehat{K}\geq K+k)$ from $\alpha^{\lfloor{k/2}\rfloor}$ to $\alpha^{k}$ for all $k\in\mathbbm{N}$.
$\mathcal{C}_{k}\coloneqq\biggl\{g=\sum_{j=0}^{k}\theta_{j}\mathbbm{1}_{{I}_{j}}\;:\;\bigsqcup_{j=0}^{k}I_{j}=[0,1),\,T_{\mathring{I}_{j}}(Z,\theta_{j})-q_{\alpha}(\lvert\mathring{I}_{j}\rvert)\leq 0,\,j=0,\dots,k\biggr\},$ (5)
where $\mathring{I}$ denotes the interior of interval $I$. Note that in this particular form of $\mathcal{C}_{k}$ we control the type I error locally on every segment $I_{i}$. Further, to avoid overfitting to the data, we minimize the complexity of candidates, which is measured by the number of change points, and define
@@
Note that $\tilde{f}=\sum_{i=1}^{n}Z_{i}\mathbbm{1}_{[x_{i},x_{i+1})}$ clearly lies in $\mathcal{C}_{n-1}$, so $\widehat{K}$ is finite and takes values in $\{0,1,\ldots,n-1\}$. Thus $\mathcal{C}_{\widehat{K}}\neq\emptyset$ and consequently a solution $\widehat{f}$ of (7) always exists. Owing to the specific structure of the optimization problem, $\widehat{f}$ can be efficiently computed in the dynamic programming framework (see Section 4).
@@ 2.3 Comparison with global error control
As a simulation study, we consider the recovery of the teeth function that contains $K=80$ change points under the additive noise model (1), where the random errors are scaled Student’s $t_{3}$-distributed with standard deviation equal to one. The simulation results are given in Figure 2 and Table 1. Clearly, MUSCLE ($\alpha=0.3$) recovers nicely the number and the locations of change points, while MQS ($\tilde{\alpha}=0.3$) severely underestimates the number of change points. To enhance the detection power of MQS, we adjust $\tilde{\alpha}=1-(1-0.3)^{81}\approx 0.999$. This adjustment significantly improves MQS in the estimation of $K$, bringing its performance closer to MUSCLE, but at the expense of nearly no statistical guarantee. Still, MQS with such a large $\tilde{\alpha}$ underperforms MUSCLE in estimating the change point locations as well as the whole signal. See Section 5.1 (and also Section 1.2) for further comparisons.
@@
Table 1: Comparison of MQS and MUSCLE in the setup of Figure 2. For each criterion, the average value and the standard deviation (in brackets), over 200 repetitions, are presented. Here $d_{H}$ is the Hausdorff distance between estimated and true change points, MISE the mean integrated squared error of $\widehat{f}$ and MIAE the mean integrated absolute error. Their formal definitions are given in (12), (23) and (24) in Appendix B, respectively.
@@
$\widehat{K}$ 80(0.4)80\;(0.4) 40(0.5)40\;(0.5) 77(1.6)77\;(1.6)
$d_{H}$ 0.001(0.0030)0.001\;(0.0030) 0.015(0.0037)0.015\;(0.0037) 0.008(0.0006)0.008\;(0.0006)
MISE 0.1(0.041)0.1\;(0.041) 2.7(0.065)2.7\;(0.065) 1.8(0.108)1.8\;(0.108)
MIAE 0.2(0.020)0.2\;(0.020) 1.4(0.008)1.4\;(0.008) 0.9(0.057)0.9\;(0.057)
@@ 3 Statistical Properties
3.1 Consistency and localization rates
@@
$F_{\min,k}=\underset{i:x_{i}\in I_{k}}{\min}F_{i}\quad\text{and}\quad F_{\max,k}=\underset{i:x_{i}\in I_{k}}{\max}F_{i}.$
For $k=1,\ldots,K$, let $\delta_{k}=\theta_{k}-\theta_{k-1}$ and refer to its absolute value as jump size. We use
$\xi_{k}=\min\left\{\xi_{F_{\min,k},\beta}(\delta_{k}/2),\xi_{F_{\max,k},\beta}(-\delta_{k}/2)\right\}$
to characterize the quantile jump at $\tau_{k}$. In case of $F_{i}$ being continuously differentiable, $\xi_{k}$ is equivalent to $\lvert\delta_{k}\rvert$ up to a constant if $\lvert\delta_{k}\rvert$ is small, see Example 1. Further, we define
$\Lambda\coloneqq\underset{k=1,\dots,K}{\min}\lvert\tau_{k}-\tau_{k-1}\rvert,\quad\Theta\coloneqq\underset{k=1,\dots,K}{\min}\xi_{k}\;\text{ and }\;\Omega\coloneqq\underset{k=1,\dots,K}{\min}\sqrt{\lambda_{k}}\xi_{k}.$ (8)
@@
    $\liminf\Lambda_{n}>0$ and $\sqrt{n}\Omega_{n}-\max_{m\leq n}q_{\alpha_{n}}(m)/\sqrt{2}\to\infty$.
@@
    $\liminf\Lambda_{n}=0$ and for some constant $c>2$,
    $\sqrt{n}\Omega_{n}\geq c\sqrt{-\log\Lambda_{n}}\;\text{ and }\;(c-2)\sqrt{-2\log\Lambda_{n}}-{\max_{m\leq n}q_{\alpha_{n}}(m)}\to\infty.$
@@
By Lemma 10 in Appendix A, we have $\max_{m\leq n}q_{\alpha_{n}}(m)\lesssim\sqrt{-\log\alpha_{n}}.$ This allows us to select $\alpha_{n}$ approaching zero at certain explicit rates to satisfy the above conditions on $q_{\alpha_{n}}(m)$. We discuss the implication of Theorem 3 under conditions (i) and (ii), separately.
@@
    Note that $\liminf_{n}\Lambda_{n}>0$ implies that $\sup K_{n}<\infty$, that is, the number of change points stays bounded. In particular, if the quantile function $f=f_{n}$ in Model 1 stays the same for different $n$, MUSCLE with any vanishing sequence $\alpha_{n}\searrow 0$ estimates consistently the number of change points.
@@
    Consider the example of $f_{n}=\Delta_{n}\mathbbm{1}_{[0.5,0.5+\Lambda_{n})}$ with additive standard Gaussian noise and jump size $\Delta_{n}>0$, which corresponds to Model 1 with $\beta=0.5$. By simple computation we have
@@
Let $T$ and $\widehat{T}$ denote the sets of true change points and estimated change points, respectively. The localization error between $T$ and $\widehat{T}$ is defined by
@@
$d_{H}\left(T,\widehat{T}\right)\coloneqq\max\left\{d(T;\widehat{T}),d(\widehat{T};T)\right\}=\max\left\{\underset{k}{\max}\,\underset{j}{\min}\,\left|\tau_{k}-\widehat{\tau}_{j}\right|,\underset{j}{\max}\,\underset{k}{\min}\,\left|\tau_{k}-\widehat{\tau}_{j}\right|\right\}.$ (12)
Theorem 4 (Localization rate in Hausdorff distance).
@@
$\mathbb{P}\left(d_{H}(T_{n},\widehat{T}_{n})\geq\varepsilon_{n}\right)\leq\alpha_{n}+4K_{n}\exp\left(-n\varepsilon_{n}\Theta_{n}^{2}\right)+4K_{n}\exp\left[-\Bigl(\sqrt{n\varepsilon_{n}}\Theta_{n}-\tfrac{\underset{m\leq n}{\max}q_{\alpha_{n}}(m)}{\sqrt{2}}-\sqrt{\log\tfrac{2e}{\varepsilon_{n}}}\Bigr)_{+}^{2}\right]$ (13)
@@
$\varepsilon_{n}=\frac{C\log n}{n\Theta_{n}^{2}}\quad\text{for some sufficiently large constant C.}$
@@ 3.2 Control of false positives
${\rm OER}\coloneqq\mathbb{E}\left[\frac{(\widehat{K}-K)_{+}}{\max\{\widehat{K},\,1\}}\right]$ (15)
@@ 4 Algorithm and Implementation
$\widehat{K}[j]\coloneqq\min\Bigg\{k\,:\underset{0\leq i\leq k}{\max}T_{\mathring{I}_{i}}(Z,\theta_{i})-q_{\alpha}(|\mathring{I}_{i}|)\leq 0,\text{for some }f=\sum_{i=0}^{k}\theta_{i}\mathbbm{1}_{I_{i}}\text{ with }\bigsqcup_{i=0}^{k}I_{i}=\bigl[0,\tfrac{j}{n}\bigr)\Bigg\}.$
@@
$\mathcal{I}\left(I\right)=\begin{cases}1&\text{if there is a $\theta$ such that }T_{\mathring{I}}(Z,\theta)\leq q_{\alpha}(\lvert\mathring{I}\rvert),\\ 0&\text{otherwise.}\end{cases}$
@@
Finally, combining the wavelet tree and the dynamic programming with pruning techniques, we obtain Algorithm 1 (in Appendix C) for computing the global optimal solution to the non-convex optimization problem defining MUSCLE.
@@
MUSCLE, defined as the global optimal solution to the optimization problem in (5)–(7), can be computed by a pruned dynamic program with the wavelet tree data structure, which has a space complexity (i.e., amount of memory) of $O(n\log n)$ and a worst-case computational complexity (i.e., runtime) of $O(n^{4}\log n)$. Further, if we use the intervals of dyadic lengths in MUSCLE, the worst-case computational complexity is $O(n^{3}\log^{2}n)$.
@@ 5.1 Simulation studies
We investigate the empirical performance of MUSCLE and its splitting-merging variant MUSCLE-S (see Remark 4) under various situations. To benchmark the performance, we include PELT (Killick et al., , 2012), WBS (Fryzlewicz, , 2014), SMUCE (Frick et al., , 2014), FDRSeg (Li et al., , 2016) and MQS (Vanegas et al., , 2022) in Gaussian models, and R-FPOP (Fearnhead and Rigaill, , 2019), KSD (Madrid Padilla et al., , 2021), MQS (Vanegas et al., , 2022), NOT-HT (Baranowski et al., , 2019, the heavy tailed version), ED-PELT (Haynes et al., , 2017), RNSP (Fryzlewicz, , 2024, the CUSUM version) in non-Gaussian models, as competitors. For implementation, we compute WBS from R package wbs, PELT from changepoints, SMUCE and FDRSeg from FDRSeg, MQS from mqs, NOT-HT from not, ED-PELT from changepoint.np, all available from CRAN. The implementations of R-FPOP, KSD and RNSP are available on Github (https://github.com/guillemr/robust-fpop; https://github.com/hernanmp/NWBS; https://github.com/pfryz/nsp). Default tuning parameters are used for PELT, R-FPOP, WBS, KSD and NOT-HT. For SMUCE, MQS, FDRSeg MUSCLE, MUSCLE-S and RNSP, we select $\alpha=0.3$. The length of pieces in MUSCLE-S is set to $300$. We use the max_sign_cusums criterion (see Sen and Srivastava, , 1975 and Fryzlewicz, , 2024) for estimating change points in RNSP.
As quantitative evaluation criteria of change points, we consider the estimated number $\widehat{K}$ of change points, the localization error $d(\cdot;\cdot)$ in (11), the Hausdorff distance $d_{H}(\cdot,\cdot)$ in (12), and the FDR in (14). We also include the standard criteria of function estimation, consisting of the mean integrated square error (MISE) and the mean integrated absolute error (MIAE), formally defined in (23) and (24) in Appendix B. Additionally, as change point problems can be viewed as clustering problems, we incorporate the V-measure (Rosenberg and Hirschberg, , 2007), which ranges from 0 to $1$ with larger values indicating higher accuracy.
@@ 5.1.1 Changes in median
We consider the additive noise model in (1) with $\beta=0.5$, i.e., noise terms $\varepsilon_{i}$ are independent random variables with median zero. The test signal in the Gaussian scenario (E1) has $K=2$ change points at $986$ and $1016$ with sample size $n=2000$. The values between change points are 4-4, 0 and $4$. In other scenarios (E2)(E5) we use the blocks signal (Donoho and Johnstone, , 1994) of length $n=2048$. Its change points are 205, 267, 308, 472, 512, 820, 902, 1332, 1557, 1598, 1659 and the values between change points are 0, $14.64$, 3.66-3.66, $7.32$, 7.32-7.32, $10.98$, 4.39-4.39, $3.29$, $19.03$, $7.68$, $15.37$, 0. The noise distributions are specified as follows.
@@
    Normal distribution $\mathcal{N}(0,0.9)$.
@@
    Scaled $t$-distribution $2^{-0.5}\cdot\bigl(8\cdot\mathbbm{1}_{\{1\leq i\leq 389\}}+0.5\cdot\mathbbm{1}_{\{390\leq i\leq 666\}}+4\cdot\mathbbm{1}_{\{667\leq i\leq 1445\}}+\mathbbm{1}_{\{1446\leq i\leq 2048\}}\bigr)\cdot t_{3}$.
@@
    Scaled Cauchy distribution $\bigl(0.6\cdot\mathbbm{1}_{\{1\leq i\leq 389\}}+0.05\cdot\mathbbm{1}_{\{390\leq i\leq 666\}}+0.6\cdot\mathbbm{1}_{\{667\leq i\leq 1445\}}+0.2\cdot\mathbbm{1}_{\{1446\leq i\leq 2048\}}\bigr)\cdot{\rm Cauchy(0,1)}.$
@@
    Scaled and centered $\chi^{2}$-distribution $6^{-0.5}\cdot\bigl(6\cdot\mathbbm{1}_{\{1\leq i\leq 389\}}+0.5\cdot\mathbbm{1}_{\{390\leq i\leq 666\}}+6\cdot\mathbbm{1}_{\{667\leq i\leq 1445\}}+2\cdot\mathbbm{1}_{\{1446\leq i\leq 2048\}}\bigr)\cdot\bigl(\chi_{3}^{2}-\text{median}(\chi_{3}^{2})\bigr).$
@@
    Mixed distribution $\mathbbm{1}_{\{1\leq i\leq 389\}}\cdot\mathcal{N}(0,64)+1/({2\sqrt{3}})\cdot\mathbbm{1}_{\{390\leq i\leq 666\}}\cdot t_{3}+{4}/{\sqrt{6}}\cdot\mathbbm{1}_{\{667\leq i\leq 1445\}}\cdot\bigl(\chi_{3}^{2}-\text{median}(\chi_{3}^{2})\bigr)+0.1\cdot\mathbbm{1}_{\{1445\leq i\leq 2048\}}\cdot\text{Cauchy(0,1)}$.
In all scenarios, the simulations are repeated 200 times. We display simulation results in scenarios (E1)(E2) and (E5) in Figures 3, 4 and 5, respectively. Further detailed results as well as scenarios (E3) and (E4) are given in Appendix B (Figures 9, 10 and 2). In the Gaussian scenario (E1), FDRSeg and WBS overestimate the number of change points, leading to a relatively large FDR. In contrast, PELT, SMUCE and MUSCLE provide more accurate estimation of $K$. Despite MUSCLE having a slightly larger MISE compared to SMUCE and FDRSeg, it outperforms PELT and WBS. Overall, the performance of MUSCLE is comparable to the procedures tailored for the Gaussian model. In non-Gaussian scenarios (E2)(E5), there are 11 change points in median, and 14 change points in distribution. We compute localization error, Hausdorff distance, FDR and V-measure with respect to the change points in median, and MISE and MIAE with respect to the median function. We discuss the performance of each method in detail as follows.
@@
  • KSD aims at the detection of distributional changes, but often fails to detect all of $14$ change points. The detected change points contain only a subset of change points in median, as revealed by Hausdorff distance and MISE (as well as localization error, MIAE and V-measure in Table 2).